OpenStudy (anonymous):

f(x)=(x+2)^2-4 vertex? minimum value is f(x)=?

OpenStudy (anonymous):

y = (x+2)^2 - 4 y +4 = (x+2)^2 (x+2)^2 = y+4 (x-h)^2 = y-k (standard form) where (h, k) is the vertex so the vertex is at (-2, -4)

OpenStudy (anonymous):

the minimum would be the vertex in this case since it is extending upward. so f(x) = -4

OpenStudy (anonymous):

(at x = -2)